consume two goods, What is Min Income required to generate TU=2000 given Tu=10xy

Py= 20 .

Px= 10
Answer #1

The minimization function would be

L = 10x + 20y +β(2000-10xy)

FOC:

X : 10-10βy=0 (1)

Y: 20-10βx =0 (2)

β: 2000- 10xy=0 (3)

From (1) and (2)

Y/x = 1/2

X= 2y

Putting in (3)

10.(2y).y =2000

y= 10

x = 20

The minimum income will be 10x +20y = 10(20)+20(10)= 400